Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
July 9, 2007

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen Williams
Announcements
  1. Purchasing is open again, but we need to be thrifty.
  2. Each department is going to have its own supply budget. Training and ordering information to follow.
  3. The release of the new version of ALEPH 18 is going to be delayed.
  4. We will not consider rolling out VISTA until ALEPH 18 is announced, in action and debugged.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders are flowing again - will be getting caught up. Should be ordering the locks for public machines today.
Liaisons
  1. Staff training for VISTA - on hold.
Networking
  1. Mobile devices still not synching with the Exchange server. They are unable to receive email. Will is working on it.
  2. Anti-virus server - has not been working correctly. It is not pushing the latest version. Logan is currently testing it and hope to have it back up this evening.
  3. Xerox - just received reports today that they are unable to print from ARES on the 2nd floor of West, but they can on the third. Also had a report of another problem right before this meeting.
  4. Systems to VISTA - could roll as an upgrade.
  5. SMS - Logan is testing a few file things, and then he'll talk to CNS.
  6. VM Ware - need to get the training scheduled. Talk to Business Services.
  7. NT2R1 - was dead on Friday. Several reboots - it still would not come up. Finally, Will did a consistency check and it corrected a problem, and then indicated it needed a restart. At that point, it came back up.
  8. UFDC development job that fails - it is waiting for a tape change. We should move the starting time back on that one. Set it to start right after the NT2R job.
  9. SAN - Logan has everything racked. He is in the process of reading through how it should work. Issue of switch return discussed.
Software Projects Status
  1. Grover - Stephen and Chris completed the mock-up on paper. They presented it to Winston and now have to make some changes. When that is complete, they will present it to the rest of the group to see how it fits into the workflow.
  2. DLC - Mark is working on a new database. As of July 15, all development of Ephemeral Cities ends. Erich has agreed to stop all new projects with Tracking and Greenstone. DLC has new furniture and has been reconfigured. Mark will send an updated list of port changes when complete.
Web Projects
  1. Friday the announcement about the ALEPH 18 upgrade was posted. It was removed today, as the implementation will be delayed.
Students
  1. New student hired: Mindy Stephenson.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page